What is the coupon rate of a 8-year, $1000 bond with coupons paid semiannually and a price of 6900, if it has a yield to maturity of 8.5%.

what is the coupon payment of a 4-year $1000 bond, 7.5%YTM, and with a 4% coupon rate and quarterly payments

subject: finance
Answer #1

First question:

The price is given as 6900 which appears to be abnormal compared to the face value of $1,000. It is assumed that the price given is $900.

Coupon rate = 6.751796% calculated as follows:

D10 =PMT(D8,D6,D9,02,0) A B C D 1 Formula Q#1 2 Face Value (F) Given FV $1,000 3 Interest frequency Given Semi annual 4 Numbe

Second question:

Coupon payment= FV*C/t

Where

FV= Face value (given as $1,000)

C= Coupon rate (given as 4%) and

T= number of times coupon is paid a year (given as 4, being quarterly)

Plugging the values,

Coupon payment= 1000*4%/4 = $10